OpenStudy (anonymous):

Explain how we evaluate an algebraic expression for given values of the variables. How would we evaluate the expression "3x - 4y + 9" when x = 2 and y = 5?

OpenStudy (bahrom7893):

substitute 2 for x and 5 for y: 3*2 - 4*5 + 9 = 6 - 20 + 9 = -5